simplify the expression -2k^2m4(-5km^3)

Found 2 solutions by checkley71, edjones:
Answer by checkley71(8405)

IF YOU MEAN TO MULTIPLY THESE 2 TERMS THEN:
-2KM^4(-5KM^3)=10K^2M^7
-----------------------------------
IF YOU MEAN TO SUBTRACT THEM THEN:
-2KM^4-5KM^3=-KM^3(2M+5)

Answer by edjones(4071)

-2k^2m4(-5km^3)
=-2k^2-5km^7
=-k(2k+5m^7) or -2k(k+2.5m^7) whichever you like.
Ed